40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1495     Entry time: Sun Apr 19 03:34:05 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Saturday night lock 
Tonight I was able to go up to arm power = 33, by mainly tweaking the DARM gain. A small progress.
In order to give more phase margin to the CARM MC_L path, I added a 300:100 filter to C1:LSC-MC.
To reduce the load to the lsc computer I deleted several filters from the filter bank, which were not used in the locking scripts.
Before I deleted the filters, I checked in the current chans directory into the svn repository.
If you want to restore the deleted filters, go back to the revision 36142.
ELOG V3.1.3-